OpenStudy (anonymous):

Find the n th term and sum of n terms of the series 5 + 8 + 17 + 44 ..........

OpenStudy (anonymous):

Ah, this is a nice series. \[[5,8,17,44]\]

OpenStudy (experimentx):

find the difference .. note the difference in the difference is a multiple of 3

OpenStudy (anonymous):

Yes the first order difference is in Gp

OpenStudy (anonymous):

and on...so do as @experimentX suggests, first examine any sort of relationship that could exist, you see that each termer is greater than the next

OpenStudy (anonymous):

I suggest that you show us the values you have for the difference between the numbers in the series

OpenStudy (anonymous):

Once, you have those differences, see if there is a pattern or relationship among them

OpenStudy (zepp):

Looks like 3 to some power :D

OpenStudy (experimentx):

tn+1 = tn + 3^(n-1)

OpenStudy (anonymous):

term1 = 5 term2 = 8 term2 - term1 = 3 so term2 = term1 + 3 term3 = 17 term3 - term2 = 9 so term3 = term2 + 9 term4 = 44 term5 - term4 = 27 so term4 = term3 + 27 term5 = ...

OpenStudy (experimentx):

don't know closed form at the moment!!

OpenStudy (anonymous):

So @Yahoo! How are you doing with this?

OpenStudy (anonymous):

Ar^n + b1n^k-1 + b2n^k-2............bk

OpenStudy (anonymous):

Hmm, can you explain the terms please?

OpenStudy (anonymous):

it is the general formula lol1

OpenStudy (anonymous):

ok look at the sequence first term is (3^1+7)/2=5 second is (3^2+7)/2=8 third is (3^3+8)/2=17 can you now tell me general formula ?

OpenStudy (experimentx):

i guess we should use summation rule